a bird species in danger of extinction has a population that is decreasing (A=A(base0)e^kt). five years ago the population was at 1400 and today only 1000 are still alive. once the population drops below 100, the situation is irreversible. When will this happen?
Answer by stanbon(48558)

a bird species in danger of extinction has a population that is decreasing (A=A(base0)e^kt). five years ago the population was at 1400 and today only 1000 are still alive. once the population drops below 100, the situation is irreversible. When will this happen?
----------------------------
A(t) = A(o) e^(kt)
Find "k":
1000 = 1400 e^(k*5)
(5/7) = e^(5k)
Take the ln to get:
5k = ln (5/7)
5k = -0.3365
k = -0.06729
---------------
Equation: A(t) = A(o) e^(-0.06729t)
----------
When will the population be 100?
100 = 1000 e^(-0.06729t)
0.10 = e^(-0.06729t)
-0.06729t = ln(0.10)
t = 31.22 years from 5 yrs. ago
OR
26.22 years from NOW
=======================
Cheers,
Stan H.